1 A force of 250.0 N is exerted on a 1,250 N car while moving it a distance of 5.00 m. How much work (in J) was done on the car? Enter answer 2. An electric hoist does 133,000 J of work in raising a 350.0 kg load. How high (in m) was the load lifted? Enter answer 3 What is the kinetic energy (in J) of a 45-gram bullet that is traveling at 185 m/s? Enter answer 4 How much work (in J) is done in raising a 65.0 kg crate a distance of 4.50 m above a storeroom floor? Enter answer 5. A 60.0 kg hiker is able to ascend a 840 m high slope in 1 hour 21 minutes. What was the average power output of the hiker in W? Enter answer
Answer #1

(1) The work is, W = Fs = (250 N)(5 m) = 1250 J (2) The energy is, E = mgh h=E mg 133,000 J (350 kg) (9.8 m/s) = 38.8 m (3) T

(4)The work done is, W = (65 kg (9.8 m/s (4.50 m = 2867J (5) The work done is, W = (60 kg)(9.8 m/s?)(840 m) = 4.94x10% ] The